OpenStudy (anonymous):

please, give me a reference about the precession perihelion of newton theory help me!

OpenStudy (anonymous):

principia mathematica lol

OpenStudy (anonymous):

you have about it?

OpenStudy (anonymous):

why so \[\frac{ d \hat{r} }{ dt } = \hat{\theta}\frac{ d \theta }{ dt }\] and \[\frac{ d \hat{\theta} }{ dt } = - \hat{r} \frac{ d \theta }{ dt }\] in pictures